OpenStudy (anonymous):

Can someone help me with four Algebra 2 questions?

OpenStudy (anonymous):

1. Write the equation of the quadratic function with roots -10 and -8 and a vertex at (-9, -3). 2. Write the equation of the quadratic function with roots 0 and 1 and a vertex at (1/2, -2/3) 3. Write the equation of the quadratic function with roots -9 and and -3 and a vertex at (-6, -1). 4. Write the equation of the quadratic function with roots -2 and 4 and a vertex at (1, 9).

OpenStudy (anonymous):

1. the quadratic equation is given by x^2-(a+b)X+ab where a,b are the roots since it is given that this passes through a point (x1,y1) your quadratic equation simply equated to y will be the answer

OpenStudy (anonymous):

same procedure is applicable to all the 4 after writing the equation just verify it by giving x the value of x1

OpenStudy (anonymous):

Would you mind telling me the answer for each one and explaining how you got it. I've got a final exam coming up and I'm trying to jot some things down.

OpenStudy (anonymous):

in the first case roots are -10 and -8 then a=-10 b=-8 a+b=-18 ab=80 the quadratic expression is x^2-(-18)x+80 just equate this to y therefore the equation is y=x^2+18x+80 given vertex is (-9,-3) implies the above equation should satisfy for the given vertex values of x and y implies x=-9 should fetch the value of y=-3 put x=-9 then y=(-9)^2+18(-9)+80 should fetch you -3 but the answer is -1 so just multiply both sides by 3 so your equation becomes y=3(x^2-18x+80)

OpenStudy (anonymous):

Thank you... But what about the others?